OpenStudy (anonymous):

g(x) =x^2 + 3x +3, what does each part stand for if y=mx+b, m=slope, x=x-axis and m=slope

OpenStudy (anonymous):

your b is 3 so when x=0 y=3 you actually have a quadratic equation so you can factor is to get the x intercepts x^2+3x+3=-b+/-sqrt(b^2-4ac)/2a then with those 3 points you can graph your quadratic :)

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

in this case; y = M(x-h)^2 + b

OpenStudy (amistre64):

g(x) =x^2 + 3x +3 g(x) =(x^2 + 3x + 9/4) +3 +9/4 g(x) =(x+3/2)^2 +21/4 M = "1", y intercept = 21/4

OpenStudy (anonymous):

sorry amistre but when x=0 g(0)=3/2^2+21/4 9/4+21/4=30/4 i think you ment to say -9/4 not +9/4 at the end of the 1st part you did because your adding "0"

OpenStudy (anonymous):

which in turn would look like this which amistre stated g(x)=(x+3/2)^2+3/4

OpenStudy (amistre64):

.... good eye :)

OpenStudy (amistre64):

im used to doing that manuever on the other side of an equals sign ;)

OpenStudy (anonymous):

standard :)